OpenStudy (jayden):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11). is the answer y=4x+75?

OpenStudy (jayden):

please someone help

OpenStudy (tkhunny):

Immediately - No. If the slope is perpendicular to y = -4x - 2, the slope must be 1/4, not 4.

OpenStudy (tkhunny):

Perpendicular slopes, as long as one line is not vertical, multiply to -1. We have -4. This requires - 4 (1/4) = -1 Finally (y+11)=(¼)(x+16)

OpenStudy (jayden):

so it is written like this -11=1/4(x--16)

OpenStudy (tkhunny):

Not quite. You're missing the 'y-' in the front.

OpenStudy (jayden):

so y--11=1/4(x--16) which comes out to be y+11=1/4(x+16) which comes out to be y+11=1/4x+4.25 am i on the right path

OpenStudy (jayden):

and the answer would be y=1/4x-6.75

OpenStudy (tkhunny):

(1/4)*16 = 4

OpenStudy (jayden):

ok i missed a step it would equal y=1/4x+15 i am sorry that i am not picking this up right away.

OpenStudy (tkhunny):

You can do it. Slow down and think it through. Your hand wants to go faster than your brain. It is not uncommon. :-) You'll get it with a little patience.

OpenStudy (jayden):

ok when i do this function 1/4 (x--16) in my calculator the answer i get is 17/4 or 4.25 is that right

OpenStudy (tkhunny):

Must be something funny about how it is interpreting the "--". Use "+" there, instead. For the future, you may want NEVER to write "--" again. Always write "+" for that. "--" is quite confusing.

OpenStudy (jayden):

ok

OpenStudy (tkhunny):

Maybe 'x' has the value '1' for some reason. And thus we see why the brain remains of more valuable than the calculator.

OpenStudy (jayden):

still get 4.25 and i did 1/4(x+16)

OpenStudy (jayden):

ow ok thank you for that comment

OpenStudy (jayden):

so the right answer would be y=1/4x+15

OpenStudy (tkhunny):

Almost. 4 - 11, not 4 + 11. Careful with your algebra. The notation should help you. Be consistent. Don't be afraid to rewrite the WHOLE equation as many times as necessary. Make sure your can follow your tracks.

OpenStudy (jayden):

no that would not be the right answer it would be y=1/4x-7 cause you subtract 11 not add 11

OpenStudy (tkhunny):

NOW we're talking! Good work.

OpenStudy (jayden):

thank you so the answer for this question, Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4) would be y=x+4

OpenStudy (jayden):

tk hunny are you still there

OpenStudy (tkhunny):

Not sure just by looking at it. What does the algebra say? (2,4) (5,4) Whoops! I notice something right off. The y-coordinates are the same! y = 4 Done!

OpenStudy (jayden):

the question is Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4) and my answer to this question is y=x+4 or is the answer y=4

OpenStudy (tkhunny):

Try the two points in your equation. Neither one works. I don't know how you got that equation, but it is not particularly related to the two given points.

OpenStudy (jayden):

\[4-4\div5-2=0/3 \] so m=0 and you would find b by doing 4=0(2)=b and you would get 4=0 so be would equal 4 so then you would put it into the formula y=mx+b and that would look like y=4 because their is no slope so their would be no x. does this look right to you

OpenStudy (jayden):

i think i just solved my own problem

OpenStudy (tkhunny):

You do have to be careful with vertica and horizontal lines. y = mx + b When m = 0, it look s a little odd. y = b. This is a horizontal line. Vertical lines are worse, x = p, where p is the x-intercept that we're not used to seeing.

OpenStudy (jayden):

ok so is y=4 the correct answer and thank you sooo much for all the help that you are giving me it helps tremendously

OpenStudy (tkhunny):

You are very welcome.

OpenStudy (jayden):

can i run another answer by you to check to see if it is right

OpenStudy (jayden):

write an equation of a line in slope intercept form that has a slope of -3/4 and passes through the point (4,5). Is the answer y=3/4+8?

OpenStudy (jayden):

my work is 5=-3/4(4)=b which would equal 5=-3 plus three on both sides and you get eight and the slope is the reciprocal of -3/4 so it would be y=4/3x+8

OpenStudy (tkhunny):

Let's see (y-5)=-¾(x-4) y - 5 = -¾x + 3 y = -¾x + 8 1) You lost your 'x'. That has to be just carelessness. Be more careful. 2) We started with -¾. How did yours switch to +¾? You're close. A little more caution.

OpenStudy (jayden):

so the answer is Y=-3/4x+8

OpenStudy (tkhunny):

the 'x' and 'y' are part of the equation. You cannot just ignore them and magically put them back in at the end.

OpenStudy (jayden):

ok

OpenStudy (jayden):

so it would be y-5=-3/4(x-4) which would then equal y=-3/4x+8

OpenStudy (jayden):

i see how you got that answer thank you

OpenStudy (tkhunny):

That's it. :-) I'm out.

OpenStudy (jayden):

ok thank you